# Hot Rocks Survey III: A deep eclipse for LHS 1140c and a new Gaussian process method to account for correlated noise in individual pixels

### v1

@@
Time-series photometry at mid-infrared wavelengths is becoming a common technique to search for atmospheres around rocky exoplanets. This method constrains the brightness temperature of the planet to determine whether heat redistribution is taking place - indicative of an atmosphere - or whether the heat is reradiated from a low albedo bare rock. By observing at 15m we are also highly sensitive to CO2 absorption. We observed three eclipses of the rocky super-Earth LHS 1140c using MIRI/Imaging with the F1500W filter. We found significant variation in the initial settling ramp for these observations and identify a potential trend between detector settling and the previous filter used by MIRI. We analysed our data using aperture photometry but also developed a novel approach which joint-fits pixel light curves directly using a shared eclipse model and a flexible multi-dimensional Gaussian process which models changes in the PSF over time. We demonstrate using simulated data that our method has the ability to weight away from particular pixels which show increased systematics, allowing for the recovery of eclipse depths in a more robust and precise way. Both methods and an independent analysis detect the eclipse at and are highly consistent with a low albedo bare rock. We recover a dayside brightness temperature of K, close to the theoretical maximum of K. We rule out a wide range of atmospheric forward models to including pure CO2 atmospheres with surface pressure mbar and pure H2O atmospheres with surface pressure bar. Our strict constraints on potential atmospheric composition, in combination with future observations of the exciting outer planet LHS 1140b, could provide a powerful benchmark to understand atmospheric escape around M dwarfs.
@@
The search for atmospheres on rocky exoplanets is one of the major science goals defining the early JWST era. However, we still lack definitive evidence of an atmosphere around any rocky planet, with only a few hints of signals reported to date (e.g. August et al. 2025; Hu et al. 2024; Patel et al. 2024; Gressier et al. 2024; Banerjee et al. 2024). At present, M-dwarf systems offer the best hope of characterising rocky exoplanet atmospheres using current facilities and they have therefore become the targets of numerous JWST programs (e.g. Greene et al. 2023; Zieba et al. 2023b; Weiner Mansfield et al. 2024; Alam et al. 2025). It is still an open question how many M-dwarf targets can retain atmospheres due to the significant flaring and intense EUV emission these stars are believed to exhibit in their first Gyr (Shields et al. 2016). There is a large degree of uncertainty around the extent of this activity however, and a planet may still retain a secondary atmosphere after this active phase finishes (Krissansen-Totton et al. 2024) or could even form a new secondary atmosphere from volcanic outgassing (Kite & Barnett 2020; Tian & Heng 2024). Ultimately the only way we will answer this question is by observing a sample of M-dwarf targets and measuring whether or not they host atmospheres.
@@
There are multiple approaches to inferring the presence of an atmosphere, with different methods varying in observational cost. We used eclipse photometry over the 15m wavelength band offered by the MIRI F1500W filter. This approach assumes that the planet being studied is tidally locked to its host star (which is expected for our target from dynamical modelling of the LHS 1140 system; Gomes & Ferraz-Mello 2020) and asserts that the surfaces most likely to exist on rocky worlds within a given temperature range (300 K ¡ ¡ 880 K) have low surface albedos of (Koll et al. 2019; Mansfield et al. 2019). As there should be insignificant heat redistribution on a planet without a substantial atmosphere (Joshi et al. 1997; Selsis et al. 2011; Koll 2022), we can expect this to result in a hot dayside which should produce strong thermal emission at mid-infrared wavelengths. Alternatively, a planet with a substantial atmosphere should display more efficient heat redistribution to its nightside, cooling the dayside and reducing its thermal emission. In addition, carbon dioxide (CO2) hosts a strong absorption band at 15m which can further reduce this eclipse depth if present.
@@
The advantages of infrared eclipse photometry are that it can require fewer observing hours to identify an atmosphere compared to observing full phase curves (Koll et al. 2019). Unlike transmission spectroscopy, our observations are not affected by the transit light source effect (TLS; Rackham et al. 2018) as we observe the planet as it passes behind its star, avoiding degeneracies in distinguishing a planetary atmosphere from stellar surface inhomogeneities. Compared with eclipse spectroscopy, photometry offers higher signal-to-noise which reduces observational cost - although at the expense of increased degeneracy in our interpretations (Hammond et al. 2025).
@@
These benefits were the motivation for the Hot Rocks Survey (GO 3730 PI: Diamond-Lowe, Co-PI: Mendonça), a large JWST Cycle 2 program to observe nine rocky planets with equilibrium temperatures ranging from K - 910 K using 15m eclipse photometry. Of these targets, LHS 1140c has the lowest equilibrium temperature (T = K; Cadieux et al. 2024a) and lowest instellation ( ; similar to Mercury at aphelion) of the sample. With a mass of and radius of it can be considered a super-Earth. This gives it a reasonably high escape velocity of km/s (using constraints from Cadieux et al. 2024a) - larger than any Solar System rocky body. Combining these properties together makes it a high priority target as the cosmic shoreline model suggests atmospheres are more likely to be retained by planets with low instellation (in particular low integrated EUV flux) and high escape velocities (Zahnle & Catling 2017). Modelling of atmospheric escape suggests LHS 1140c may well have retained a secondary atmosphere (Kite & Barnett 2020; Chatterjee & Pierrehumbert 2024), and observations from Swift (Gehrels et al. 2004) have shown the LHS 1140 system shows relatively low levels of present day NUV flux and estimated FUV flux for a star of this type (Spinelli et al. 2019). However, we note recent work suggests mid-type M dwarfs such as LHS 1140 may have had longer early active periods than previous models predicted, making atmospheric retention more challenging (Pass et al. 2025).
@@
LHS 1140c was originally discovered by Ment et al. (2019) and is the inner planet of the system. The only other confirmed planet in the system - LHS 1140b - lies within its star’s habitable zone and has a low-density with a mass of and radius of , potentially consistent with a water world or mini-Neptune (Cadieux et al. 2024a). Transit observations using NIRISS/SOSS (Cadieux et al. 2024b) and NIRSpec (Damiano et al. 2024) suggest that LHS 1140b is inconsistent with the mini-Neptune case due to the flat transmission spectrum recovered. This implies that it could be a water world with a high mean molecular weight atmosphere. The NIRISS/SOSS observations happened to capture a transit of our target LHS 1140c entirely within a transit of LHS 1140b, although there was low signal-to-noise from just a single transit and the transmission spectrum was consistent with a flat line with no molecular features or scattering-slope from hazes detected. Additionally, it has previously been noted that there is a discrepancy between the transit depths of LHS 1140c from TESS observations and a single Spitzer observation at 4.5m - suggesting possible absorption from CO2 (Cadieux et al. 2024a). However, these recent NIRISS/SOSS observations appear to show that the TESS observations underestimated the transit depth, with the new observations lying between the TESS and Spitzer transit depths. Eclipse observations using the F1500W filter can further resolve this discrepancy as our observations are highly sensitive to the presence of CO2.
@@
One of the lessons we are learning is that systematics affect at least some of the observations in our survey (e.g. August et al. 2025; Meier Valdés et al. 2025). In addition, there can be significant variation in the strength and sign of the initial settling ramp which covers at least the first minutes of our observations. We present a possible explanation for these variations related to the previous filter used by MIRI, as we note it remains in place throughout target acquisition. In addition, we present a new Gaussian Process (GP) based approach which fits the eclipse depth at the level of individual pixels without performing an aperture extraction. We also retrieve information about how the point spread function (PSF) changes in time and characterise the strength of interpixel capacitance. Our method has the benefit of being able to identify and weight away from systematics found in individual pixels, which we demonstrate using simulated data. In one of our observations, this method identifies a possible persistence effect caused by a cosmic ray.
@@
Our work is laid out as follows: first we introduce our new pixel-based approach to account for systematics in Section 2 and will demonstrate the advantages of this method using simulated data in Section 3. Those primarily interested in the observations may move straight to Section 4 which presents the analysis of our observations using the standard approach of aperture photometry as well as using our pixel-based approach. It also includes our findings on MIRI detector settling. Section 5 will describe our modelling of LHS 1140c and the interpretation of our recovered eclipse depth. Section 6 will discuss the implications of our findings and elaborate on the strengths and weaknesses of our new method and where it may be extended. Our conclusions are outlined in Section 7.
@@ 2.1 Motivation for a new pixel-fitting method
There have been a number of methods developed to analyse time-series photometric observations. After data cleaning, a standard approach is to define an aperture region where flux from the target star is greatest and add up the flux within this region with a uniform weighting on each pixel. This classic approach is known as aperture extraction and can perform well for bright targets where any remaining background noise from imperfections in the background subtraction are minimal. It has a couple of limitations, the choice of the aperture region and shape is quite arbitrary, but choosing too wide an aperture reduces signal-to-noise due to increased background noise while too small an aperture reduces signal-to-noise from missed flux. Typically an intermediate aperture radius is chosen which maximises signal-to-noise. However, applying a uniform weighting for each pixel is generally not optimal as brighter pixels will be less impacted by background noise than outer pixels and might ideally be weighted more.
@@
Optimal extraction attempts to remedy these issues by fitting for the PSF shape and estimating the Poisson noise in each pixel (Horne 1986). This makes it possible to weight each pixel in a way which maximises signal-to-noise. For observations which contain a lot of background noise, optimal extraction may result in a significant improvement in signal-to-noise. This is because optimal extraction weights each pixel based on it’s particular signal-to-noise value and can be mathematically shown to maximise signal-to-noise in the presence of uncorrelated noise.
@@
However, the situation is more complicated in the presence of time-correlated noise. The Spitzer Space Telescope was known to suffer from systematics due to ”intrapixel sensitivity” - where the efficiency of each pixel varies across the surface of the pixel. In combination with instability in the telescope pointing and undersampling of the PSF, this resulted in systematics which could be different in each pixel (Ingalls et al. 2012). The first work to measure an exoplanet eclipse depth used detrending of the PSF centroid position to correct for this effect (Charbonneau et al. 2005). The technique of BiLinearly-Interpolated Subpixel Sensitivity (BLISS) mapping was later developed to account for this, directly mapping the sensitivity across each pixel’s surface (Stevenson et al. 2012). In addition, pixel-level decorrelation (PLD) also corrected for this issue by detrending the systematics in the extracted light curve against normalised pixel light curves (Deming et al. 2015). These techniques were used in early work characterising rocky exoplanet heat redistribution using phase curves from Spitzer (Demory et al. 2016; Kreidberg et al. 2019). Other relevant techniques applied to Spitzer photometric time-series include GPs which use centroid positions as input variables (Evans et al. 2015). Independent Component Analysis (ICA) - introduced to exoplanet time-series in Waldmann (2012); Waldmann et al. (2013) - has been applied to disentangle astrophysical and systematic effects and works by calculating statistically independent trends in pixel time-series (Morello et al. 2014; Morello 2015; Morello et al. 2016). A separate approach used to account for systematics in Kepler time-series is the ”Causal pixel model”, which can distinguish systematics in each pixel by comparing them to multiple other stars on the detector (Wang et al. 2016). This approach isn’t applicable to our observations of a single star however.
@@
Time-series photometry with the MIRI/Imaging mode is expected to suffer significantly less from intrapixel sensitivity variations compared to Spitzer due to the PSF being spread out over more pixels and the telescope pointing being more stable. There is also a difference in detector technology, the Si:As detectors on Spitzer were not significantly affected by intrapixel sensitivity while the InSb detectors were (Knutson et al. 2008), MIRI uses Si:As detectors. We performed some simple tests simulating a PSF with a similar width to F1500W observations and modeling the pointing stability of JWST which confirmed that effects such as intrapixel sensitivity and errors in flat fielding should be completely negligible compared to the photometric precision of these observations - even when significant intrapixel or interpixel sensitivity variations were modelled (see Appendix B for more details).
@@
However, at our early stage in understanding systematics from JWST instruments we should be careful in assuming our observations are completely free from systematics and the more approaches we can develop to analyse these data, the less sensitive we are to the limitations of any individual method. With that in mind, we introduce our new pixel-fitting approach to analyse these data, which is radically different from existing approaches such as PLD or BLISS mapping. This method combines the benefits of optimal extraction by fitting for the shape of the PSF and accounting for the amplitude of Poisson noise in each pixel. However, it also fits for the amplitude of correlated noise in each pixel. This allows it to weight away from pixels which suffer from increased systematics, in addition to weighting based on reduced Poisson noise. This method is the first we are aware of which applies Gaussian processes to joint-fit pixel time-series in astronomical data. Our approach can be more precise and robust when systematics are present in single pixels and can be used as an alternative to aperture photometry or optimal extraction for eclipse depth measurements. It also quantifies multiple detector effects in the process, such as constraining the amplitude of background noise and the amplitude of correlated Poisson noise resulting from interpixel capacitance.

### v2

@@
Time-series photometry at mid-infrared wavelengths is becoming a common technique to search for atmospheres around rocky exoplanets. This method constrains the brightness temperature of the planet to determine whether heat redistribution is taking place, which would be indicative of the presence of an atmosphere, or whether the heat is reradiated from a low-albedo bare rock. By observing at 15m, we are also highly sensitive to CO2 absorption. We observed three eclipses of the rocky super-Earth LHS 1140c, using MIRI/Imaging with the F1500W filter. We found a significant variation in the initial settling ramp for these observations and identified a potential trend between the detector settling and the previous filter used by MIRI. We analysed our data using aperture photometry, however, we also developed a novel approach, which performs a joint fit of the pixel light curves using a shared eclipse model and a flexible multi-dimensional Gaussian process which can model changes in the PSF over time. Using simulated data, we demonstrate that our method has the ability to weight away from particular pixels that exhibit increased systematics, allowing for the recovery of eclipse depths in a more robust and precise way. Both methods, as well as an independent analysis, have detected the eclipse at , while recovering an eclipse depth consistent with a low-albedo bare rock. We measured a dayside brightness temperature of K, close to the theoretical maximum of K. We rule out a wide range of atmospheric forward models to , including pure CO2 atmospheres with surface pressure mbar and pure H2O atmospheres with surface pressure bar. Our strict constraints on potential atmospheric composition, in combination with future observations of the exciting outer planet LHS 1140b, could provide a powerful benchmark for understanding atmospheric escape around M dwarfs.
@@
The search for atmospheres on rocky exoplanets is one of the major science goals defining the early JWST era. However, we still lack any definitive evidence of an atmosphere around any rocky planet, with only a few hints of signals reported to date (e.g. August et al. 2025; Hu et al. 2024; Patel et al. 2024; Gressier et al. 2024; Banerjee et al. 2024). At present, M-dwarf systems offer the best hope of characterising rocky exoplanet atmospheres using current facilities and they have therefore become the targets of numerous JWST programs (e.g. Greene et al. 2023; Zieba et al. 2023b; Weiner Mansfield et al. 2024; Alam et al. 2025). It remains an open question as to how many M-dwarf targets can retain atmospheres due to the significant flaring and intense EUV emission these stars are believed to exhibit in their first Gyr (Shields et al. 2016). However, there is a large degree of uncertainty around the extent of this activity, additionally, a planet might retain a secondary atmosphere after this active phase finishes (Krissansen-Totton et al. 2024) or could even form a new secondary atmosphere from volcanic outgassing (Kite & Barnett 2020; Tian & Heng 2024). Ultimately, the only way we can answer this question is by observing a sample of M-dwarf targets and measuring whether or not they host atmospheres.
@@
There are multiple approaches to inferring the presence of an atmosphere, with different methods varying in observational cost. We used eclipse photometry over the 15m wavelength band offered by the MIRI F1500W filter. This approach assumes that the planet being studied is tidally locked to its host star (which is expected for our target from dynamical modelling of the LHS 1140 system; Gomes & Ferraz-Mello 2020) and asserts that the surfaces most likely to exist on rocky worlds within a given temperature range (300 K ¡ ¡ 880 K) would possess low surface albedos of (Koll et al. 2019; Mansfield et al. 2019). As there should be insignificant heat redistribution on a planet without a substantial atmosphere (Joshi et al. 1997; Selsis et al. 2011; Koll 2022), we can expect this to result in a hot dayside, which would then produce strong thermal emission at mid-infrared (MIR) wavelengths. Alternatively, a planet with a substantial atmosphere should display more efficient heat redistribution to its nightside, cooling the dayside and reducing its thermal emission. In addition, carbon dioxide (CO2) hosts a strong absorption band at 15m, which can further reduce this eclipse depth if present.
@@
The advantages of infrared eclipse photometry are that it can require fewer observing hours to identify an atmosphere compared to observing full phase curves (Koll et al. 2019). Unlike transmission spectroscopy, our observations are not affected by the transit light source effect (TLS; Rackham et al. 2018) as we observe the planet as it passes behind its star, avoiding degeneracies in distinguishing a planetary atmosphere from stellar surface inhomogeneities. Compared with eclipse spectroscopy, photometry offers a higher signal-to-noise ratio (S/N), which helps reduce the observational cost – albeit at the expense of increased degeneracy in our interpretations (Hammond et al. 2025).
@@
These benefits are the motivation behind the Hot Rocks Survey (GO 3730 PI: Diamond-Lowe, Co-PI: Mendonça), a large JWST Cycle 2 programme aimed at observing nine rocky planets with equilibrium temperatures ranging from K - 910 K, using 15m eclipse photometry. Of these targets, LHS 1140c has the lowest equilibrium temperature (T = K; Cadieux et al. 2024a) and lowest instellation ( ; similar to Mercury at aphelion) of the sample. With a mass of and radius of it can be considered a super-Earth. This gives it a reasonably high escape velocity of km/s (using constraints from Cadieux et al. 2024a), which is larger than any Solar System rocky body. Combining these properties together makes it a high priority target as the cosmic shoreline model suggests atmospheres are more likely to be retained by planets with low instellation (in particular low integrated EUV flux) and high escape velocities (Zahnle & Catling 2017). Results from models of atmospheric escape suggests LHS 1140c might well have retained a secondary atmosphere (Kite & Barnett 2020; Chatterjee & Pierrehumbert 2024) and observations from Swift (Gehrels et al. 2004) have shown the LHS 1140 system shows relatively low levels of present day NUV flux and estimated FUV flux for a star of this type (Spinelli et al. 2019). However, we note recent work suggests mid-type M dwarfs such as LHS 1140 may have had longer early active periods than previous models predicted, making atmospheric retention more challenging (Pass et al. 2025).
@@
LHS 1140c was originally discovered by Ment et al. (2019) and is the innermost known planet of the system. The only other confirmed planet in the system, LHS 1140b, lies within its star’s habitable zone and has a low-density with a mass of and radius of , potentially consistent with a water world or mini-Neptune (Cadieux et al. 2024a). Transit observations using NIRISS/SOSS (Cadieux et al. 2024b) and NIRSpec (Damiano et al. 2024) suggest that LHS 1140b is inconsistent with the mini-Neptune case due to the flat transmission spectrum recovered. This implies that it could be a water world with a high mean molecular weight atmosphere. The NIRISS/SOSS observations happened to capture a transit of our target LHS 1140c entirely within a transit of LHS 1140b, although there was a low S/N coming from just a single transit and the transmission spectrum was consistent with a flat line with no molecular features or scattering-slope from hazes detected. Additionally, it has previously been noted that there is a discrepancy between the transit depths of LHS 1140c from TESS observations and a single Spitzer observation at 4.5m. This result suggests possible absorption from CO2 (Cadieux et al. 2024a). However, these recent NIRISS/SOSS observations appear to show that the TESS observations underestimated the transit depth, with the new observations lying between the TESS and Spitzer transit depths. Eclipse observations using the F1500W filter can further resolve this discrepancy as our observations are highly sensitive to the presence of CO2.
@@
One of the lessons we have learned is that systematics affect at least some of the observations in our survey (e.g. August et al. 2025; Meier Valdés et al. 2025). In addition, there can be significant variation in the strength and sign of the initial settling ramp which covers at least the first minutes of our observations. We present a possible explanation for these variations related to the previous filter used by MIRI, as we note it remains in place throughout target acquisition. In addition, we present a new Gaussian process (GP) based approach which fits the eclipse depth at the level of individual pixels without performing an aperture extraction. We also retrieve information about how the point spread function (PSF) changes in time and characterise the strength of interpixel capacitance. Our method has the benefit of being able to identify and weight away from systematics found in individual pixels, which we demonstrate using simulated data. In one of our observations, this method identifies a possible persistence effect caused by a cosmic ray.
@@
Our work is laid out as follows: first we introduce our new pixel-based approach to account for systematics in Section 2 and we demonstrate the advantages of this method using simulated data in Section 3. Those primarily interested in the observations may move straight to Section 4 which presents the analysis of our observations using the standard approach of aperture photometry as well as using our pixel-based approach. It also includes our findings on MIRI detector settling. Section 5 describes our modelling of LHS 1140c and the interpretation of our recovered eclipse depth. Section 6 presents a discussion of the implications of our findings and elaborate on the strengths and weaknesses of our new method and where it may be extended. Our conclusions are outlined in Section 7.
@@ 2.1 Motivation for a new pixel-fitting method
There have been a number of methods developed to analyse time-series photometric observations. After the data cleaning, a standard approach is to define an aperture region where flux from the target star is greatest and add up the flux within this region with a uniform weighting on each pixel. This classic approach is known as aperture extraction and can perform well for bright targets where any remaining background noise from imperfections in the background subtraction is minimal. It has a couple of limitations, the choice of the aperture region and shape is quite arbitrary, but choosing too wide an aperture reduces the S/N due to increased background noise while too small an aperture reduces S/N from missed flux. Typically an intermediate aperture radius is chosen which maximises S/N. However, applying a uniform weighting for each pixel is generally not optimal as brighter pixels will be less impacted by background noise than outer pixels and might ideally be more strongly weighted.
@@
Optimal extraction attempts to remedy these issues by fitting for the PSF shape and estimating the Poisson noise in each pixel (Horne 1986). This makes it possible to weight each pixel in a way which maximises S/N. For observations which contain a lot of background noise, optimal extraction may result in a significant improvement in the S/N. This is because optimal extraction weights each pixel based on it’s particular S/N value and can be mathematically shown to maximise the S/N in the presence of uncorrelated noise.
@@
However, the situation is more complicated in the presence of time-correlated noise. The Spitzer Space Telescope was known to suffer from systematics due to ‘intrapixel sensitivity’, where the efficiency of each pixel varies across the surface of the pixel. In combination with instability in the telescope pointing and undersampling of the PSF, this resulted in systematics that could be different in each pixel (Ingalls et al. 2012). The first work to measure an exoplanet eclipse depth used detrending of the PSF centroid position to correct for this effect (Charbonneau et al. 2005). The technique of BiLinearly-Interpolated Subpixel Sensitivity (BLISS) mapping was later developed to account for this, directly mapping the sensitivity across each pixel’s surface (Stevenson et al. 2012). In addition, pixel-level decorrelation (PLD) also corrected for this issue by detrending the systematics in the extracted light curve against normalised pixel light curves (Deming et al. 2015). These techniques were used in early work characterising rocky exoplanet heat redistribution using phase curves from Spitzer (Demory et al. 2016; Kreidberg et al. 2019). Other relevant techniques applied to Spitzer photometric time series include GPs which use centroid positions as input variables (Evans et al. 2015). Independent component analysis (ICA) – introduced to exoplanet time series in Waldmann (2012); Waldmann et al. (2013) – has been applied to disentangle astrophysical and systematic effects and works by calculating statistically independent trends in the pixel time series (Morello et al. 2014; Morello 2015; Morello et al. 2016). A separate approach used to account for systematics in Kepler time series is the ‘causal pixel model’, which can distinguish systematics in each pixel by comparing them to multiple other stars on the detector (Wang et al. 2016). However, this approach is not applicable to our observations of a single star.
@@
Time-series photometry with the MIRI/Imaging mode is expected to suffer significantly less from intrapixel sensitivity variations compared to Spitzer due to the PSF being spread out over more pixels and the telescope pointing being more stable. There is also a difference in detector technology, the Si:As detectors on Spitzer were not significantly affected by intrapixel sensitivity while the InSb detectors were (Knutson et al. 2008), MIRI uses Si:As detectors. We performed some simple tests simulating a PSF with a similar width to F1500W observations and modeling the pointing stability of JWST, which confirmed that effects such as intrapixel sensitivity and errors in flat fielding should be completely negligible compared to the photometric precision of these observations – even when significant intrapixel or interpixel sensitivity variations were modelled (see Appendix B for more details).
@@
However, at our early stage in understanding systematics from JWST instruments, we should be careful in assuming our observations are completely free from systematics. In addition, the more approaches we can develop to analyse these data, the less sensitive we are to the limitations of any individual method. With that in mind, we introduce our new pixel-fitting approach to analyse these data, which is radically different from existing approaches such as PLD or BLISS mapping. This method combines the benefits of optimal extraction by fitting for the shape of the PSF and accounting for the amplitude of Poisson noise in each pixel. However, it also fits for the amplitude of correlated noise in each pixel. This allows it to weight away from pixels that suffer from increased systematics, in addition to weighting based on reduced Poisson noise. This method is the first we are aware of which applies GPs to joint-fit pixel time series in astronomical data. Our approach can be more precise and robust when systematics are present in single pixels and can be used as an alternative to aperture photometry or optimal extraction for eclipse depth measurements. It also quantifies multiple detector effects in the process, such as constraining the amplitude of background noise and the amplitude of correlated Poisson noise resulting from interpixel capacitance.